A chemist places 0.300 g NaNO3 in a flask and adds water until the total volume is 0.300 L.
Calculate the molarity of the solution.

concentration:

M

What is the concentration in millimolar (mM)?

concentration:
Answer #1

1)

Molar mass of NaNO3,

MM = 1*MM(Na) + 1*MM(N) + 3*MM(O)

= 1*22.99 + 1*14.01 + 3*16.0

= 85 g/mol

mass(NaNO3)= 0.300 g

use:

number of mol of NaNO3,

n = mass of NaNO3/molar mass of NaNO3

=(0.3 g)/(85 g/mol)

= 3.529*10^-3 mol

volume , V = 0.3 L

use:

Molarity,

M = number of mol / volume in L

= 3.529*10^-3/0.3

= 1.18*10^-2 M

Answer: 1.18*10^-2 M

2)

M = 1.18*10^-2 M

= 1.18*10^-2 * 10^3 mM

= 11.8 mM

Answer: 11.8 mM